Statistics students participated in an experiment to test their ability to determine when 1 minute has passed. The results are given below in seconds Find the range variance and standard deviation for the given sample data. Identify one reason why the standard deviation from this sample might not be a good estimate of the standard deviation population of adults
50 53 61 66 56
range= sec

Answer by CubeyThePenguin(3113)

range = (highest) - (lowest) = 66 - 50 = 16 sec
mean = (sum of all items)/(number of items) = 57.2 sec

standard deviation = square root of the squares of the distances of the elements to the mean

= sqrt(1/5 * (50 - 57.2)^2 + (53 - 57.2)^2 + (61 - 57.2)^2 + (66 - 57.2)^2 + (56 - 57.2)^2)
= sqrt(32.56)
= 5.706

The standard deviation may not be a good estimate because the sample size is too small to accurately represent all adults.